OpenStudy (anonymous):

Solve: 3(x − 1) − (x + 5) = 2 5 0 2 1

OpenStudy (slaaibak):

use the distributive property and solve it.

OpenStudy (anonymous):

Multilply out and then simplify. 5

OpenStudy (anonymous):

i got -2?

OpenStudy (anonymous):

\[3(x-1)-(x+5)=2 \]\[3 x-3-x-5=2 \]\[3 x-x-3-5=2 \]\[-8+2 x=2 \]\[2 x=10 \]

OpenStudy (anonymous):

Can you finish from here?